OpenStudy (fantalux):

The perimeter of a rectangle is 90 ft. Find the dimensions of the rectangle if the ratio of the length to the width is 8: 7. Which of the following would be the best equation to use to solve this problem?

OpenStudy (fantalux):

8x+7x=90

OpenStudy (fantalux):

2 (8x) + 2 (7x) =90

OpenStudy (fantalux):

8x/7x= 90

OpenStudy (fantalux):

2x + 2x=90

OpenStudy (fantalux):

Those are the possibilities.

OpenStudy (fantalux):

@KamiBug

OpenStudy (fantalux):

...

OpenStudy (fantalux):

@Vincent-Lyon.Fr

OpenStudy (studygurl14):

the second one i think

OpenStudy (fantalux):

ty

OpenStudy (studygurl14):

You're welcome.

OpenStudy (studygurl14):

remember that perimeter = 2(width)+2(length)